# The Role of Social Learning and Collective Norm Formation in Fostering Cooperation in LLM Multi-Agent Systems

### v1

@@ 3.1.5. How does this operationalize cultural evolution?
In rule-based populations, payoff-biased imitation drives high-payoff strategies to spread, with small mutations preserving exploration. In LLM populations, adaptation arises from in-context updates and stochastic decoding, so the emergence of group-beneficial norms depends on model inductive biases, decoding settings, prompt design, and retention fidelity, alongside the vote as an additional selector.
@@ 3.2. Measures of success
Following Piatti et al. (2024), we evaluate two key metrics:
@@
where $H_{\text{opt}}$ is the optimal per-round harvest that keeps the resource stock at its maximum sustainable level, which is determined by $K$ and $r$. When $\eta\left(t\right)=1$, the agents harvest at the optimal level, while $\eta\left(t\right)>1$ indicates that the agents harvest more, leading to a collapse.
@@ 4.1. Validating the Framework Design
So far, we have presented the design of the framework. In this section, we establish its effectiveness by testing well-documented hypotheses about cooperation in human societies using Agent-Based Modeling (ABM). We validate the framework along three axes: (a) punishment sustains cooperation, but if removed, cooperation declines (Shutters, 2012; Szekely et al., 2021); (b) cooperation outcomes vary with punishment strength and environmental growth rate; and (c) populations with different levels of altruism, defined by their harvest thresholds, show distinct survival patterns. All simulations are run with 10 agents. See Table 3 in the Appendix for the full list of parameters.
Figure 2 shows that once punishment is disabled (Step 15), cooperation quickly collapses and resources are depleted, confirming punishment as a key mechanism for sustaining cooperation (Ostrom, 1990). To probe the ecological dimension, we sweep punishment strength $\beta$ and growth rate $r$, finding a non-linear interaction between the two (Figure 3) that creates complex conditions where adaptive cooperation must emerge to sustain the commons. Finally, we initialize altruistic and selfish agents with distinct parameter ranges and compare all-altruist, all-selfish, and mixed populations across harsh ($r=0.2$) and rich ($r=0.6$) environments. As shown in Figure 4 in the appendix, altruistic groups perform better in harsh environments by sustaining resources, while selfish groups do better in rich environments by avoiding death from under-harvesting. Mixed groups perform best in rich environments, as the variation helps them efficiently converge toward beneficial collective norms.
@@ 4.2. LLM-Agent Simulations
Having established baseline dynamics with rule-based agents under altruistic, mixed, and selfish compositions, we now evaluate an artificial society of LLM agents initialized via context to be altruistic or selfish and ask whether cooperative norms emerge. Each action in the CPR framework is implemented with a dedicated prompt: deciding effort (Figure 9), selecting a target for punishment (Figure 10), updating one’s individual norm and proposing a collective norm (Figure 11), and voting on the community norm (Figure 12).
Agents’ initial individual norms are drawn from a small bank of short templates, conditional on type, for example, “Preserve the lake for future generations” (altruistic) and “Maximize your catch while the fish are abundant” (selfish); see Table 2 for the full set. Each agent is assigned one template at random given its type, and thereafter all decisions are made in-context from the evolving social information and the currently adopted norm.
@@ 4.2.1. Cooperation in harsh environment
In the ABM baseline, altruistic populations sustain the stock longer under harsh growth, whereas selfish populations tend to overharvest and crash. Turning to LLMs to understand whether they evolve group-benefecial norms, we observe the same pattern for larger models (claude-sonnet-4, deepseek-r1, gpt-4o): altruistic initializations survive longer (Figure 5). However, smaller models collapse early regardless of initialization; efficiency traces (Figure 14, left) show early overuse followed by rapid stock collapse. The result of ANOVA (Table 1) also supports this observation; while the performance among models significantly differed regardless of the initializations, there was no consistent trend across models driven by the altruistic ratio. Instead, the difference in the altruistic ratio showed a significant interaction effect with models, suggesting that the effect of initialization bifurcated between larger and smaller models.
@@ 4.2.2. Cooperation in rich environment
In the ABM baseline, mixed populations typically perform best in rich settings because mixed populations start from a higher variance, allowing for more efficient selection towards the optimal behaviors and norms. For LLM societies, behavior differs: with more time to adapt, smaller models often survive longer when initialized selfish, while altruistic initializations sometimes underharvest and starve (Appendix Figure 6). The absence of explicit strategy copying and reliance on in-context updates make behavior stickier to the initial norm, which explains why the mixed population is not consistently best. Larger models exhibit distinct behaviors: deepseek-r1 adapts and explores (surviving near the 50-step cap), whereas gpt-4o and claude-sonnet-4 stabilize earlier with more conservative norms (Figure 14, right; Table 4). The post-hoc test also corroborated that deepseek-r1 exhibited a significantly longer survival time compared to all other models.
@@ 4.2.3. Model-specific patterns
claude-sonnet-4 and gpt-4o typically plateau near 30 rounds, largely independent of the initial norm, whereas deepseek-r1 often reaches the 50-round cap, especially from selfish starts (Figure 6). Efficiency trajectories corroborate this: deepseek-r1 stabilizes by steps 15–20 and then nudges upward, while claude-sonnet-4 and gpt-4o settle at lower efficiency levels and remain there (Figure 14, right). The language of proposed group norms mirrors these dynamics (Table 4): deepseek-r1 quickly adjusts target effort and, after step 40, cautiously raises it; gpt-4o keeps effort targets essentially unchanged. Under identical environmental dynamics, this points to a stronger exploratory bias in deepseek-r1 and a more conservative/altruistic bias in claude-sonnet-4 and gpt-4o.
@@ 4.2.5. Ablation study: What drives cooperation?
Specifically, we compare three reduced variants against the full model: (A) Only Social Learning (OSL): agents imitate higher-payoff peers but no group norm is shared; (B) Only Group Decision (OGD): agents vote on a common norm but cannot imitate peers; and (C) Neither: both channels are removed, so agents act based only on their individual history and environmental feedback. All other parameters match the main simulations. Survival time (over n=10n\!=\!10 trials per condition) is shown in Figure 8 and Figure 13.
@@
When both channels are removed (Neither), societies consistently show the lowest survival times ($\overline{T_{s}}_{\text{Neither}}=16.22$) across environments and priors ($\overline{T_{s}}=20.98,t\left(898\right)=-2.78,p=0.006$), confirming that some form of alignment, implicit or explicit, is necessary to sustain cooperation. This establishes that coordination mechanisms, rather than individual adaptation alone, are key to stability.
@@
Conversely, retaining imitation but removing the explicit norm (OSL) consistently harms cooperation ($\overline{T_{s}}_{\text{OSL}}=17.56,\overline{T_{s}}=20.98,t\left(898\right)=-1.96,p=0.050$). Without a shared policy to anchor expectations, agents imitate whichever peers happen to achieve short-term payoffs, amplifying stochastic fluctuations and hastening collapse. The finding reveals that implicit imitation, in isolation, cannot guarantee stable group outcomes.
@@
Figure 8. Survival time comparison of deepseek-r1 gpt-4o in ablation conditions (See Figure 13 for qwen3-32b). We compared the survival time (with $\pm 1$ s.e.m.) of four conditions (All, OSL, OGD, Neither) across different priors of populations (selfish, mixed, altruistic) in harsh and rich environments. Detailed observations are discussed in Section 4.2.5.
@@ 5. Discussion & Conclusion
This paper introduced a CPR simulation framework grounded in Ostrom’s institutional design principles and cultural evolutionary theory, enabling LLM societies to develop group-beneficial norms endogenously without explicit reward signals. Through both ABM and LLM simulations, we demonstrated the validity of the framework design and its ability to elicit diverse cooperative behaviours and norms across different LLM models. The ablation results show that removing both alignment channels, social learning and group norms, consistently leads to rapid collapse across all environments and priors. This confirms that some form of coordination, whether implicit imitation or explicit norm sharing, is essential for sustaining cooperation in LLM societies. Our results establish the framework as a theoretically driven and ecologically valid testbed for studying norm evolution and cooperative dynamics in agentic societies.
@@ 5.3. Ethical considerations
Acknowledgements.
We are grateful to Levin Brinkmann for insightful discussions and valuable suggestions. This work was supported in part by the Japan Science and Technology Agency, PRESTO Grant Number JPMJPR246B.
Appendix A Prompts and additional figures
@@ 2. Propose what the others should do in the community
Figure 11. Prompt for updating the individual norm and proposing the community norm
@@
Table 2. Templates for initialising the individual norm

### v2

@@ 3.1.5. How does this operationalize cultural evolution?
In rule-based populations, payoff-biased imitation drives high-payoff strategies to spread, with small mutations preserving exploration. In LLM populations, adaptation arises from in-context updates and stochastic decoding, so the emergence of group-beneficial norms depends on model inductive biases, decoding settings, prompt design, and retention fidelity, alongside the vote.
@@ 3.2. Measures of success
Following piatti2024govsim, we evaluate two key metrics:
@@
where $H_{\text{opt}}$ is the optimal per-round harvest that keeps the resource stock at its maximum sustainable level, determined by $K$ and $r$. When $\eta\left(t\right)=1$, the agents harvest at the optimal level, while $\eta\left(t\right)>1$ indicates that the agents harvest more, leading to a collapse.
@@ 4.1. Validating the Framework Design
So far, we have presented the design of the framework. In this section, we establish its effectiveness by testing well-documented hypotheses about cooperation in human societies using Agent-Based Modeling (ABM). We validate the framework along three axes: (a) punishment sustains cooperation, but if removed, cooperation declines shutters2012punishment; szekely2021evidence; (b) cooperation outcomes vary with punishment strength gibson2005local and environmental growth rate; and (c) populations with different levels of altruism barclay2004trustworthiness, defined by their harvest thresholds, show distinct survival patterns. All simulations are run with 10 agents. See Table 3 in the Appendix for the full list of parameters.
Figure 2 shows that once punishment is disabled (Step 15), cooperation collapses faster and resources are depleted in a fater rate, confirming punishment as a useful mechanism for sustaining cooperation ostrom1990governing. To probe the ecological dimension, we sweep punishment strength $\beta$ and growth rate $r$, finding a non-linear interaction between the two (Fig. 3) that creates complex conditions where adaptive cooperation must emerge to sustain the commons. Finally, we initialize altruistic and selfish agents with distinct parameter ranges and compare all-altruist, all-selfish, and mixed populations across harsh ($r=0.2$) and rich ($r=0.6$) environments. As shown in Fig. 12 in the Appendix, altruistic groups perform better in harsh environments by sustaining resources, while selfish groups do better in rich environments by avoiding death from under-harvesting. Mixed groups perform best in rich environments, as the variation helps them efficiently converge toward beneficial collective norms. Under weaker penalties, over-harvesting is less immediately costly, so behavior can appear stable early and only diverge once cumulative stock depletion makes consequences salient.
@@
Figure 4. Altruistic groups do better in harsh environments and selfish groups do better in rich environments. We set up altruistic and selfish agents by initializing them with parameters drawn from different ranges (all in the initial range of a general agent). Then we contrast the survival time of a population of all altruists, one of all selfish agents, and one of half altruistic, half selfish agents. We ran each condition 100 times and plotted the mean and standard error. The results suggest that the altruistic population outperforms other populations in a harsh environment, while a mixed population has a better group outcome in a rich environment.
@@ 4.2. LLM-Agent Simulations
Having established baseline dynamics with rule-based agents under altruistic, mixed, and selfish compositions, we now evaluate an artificial society of LLM agents initialized via context to be altruistic or selfish and ask whether cooperative norms emerge. Each action in the CPR framework is implemented with a dedicated prompt: deciding effort (Fig. 8), selecting a target for punishment (Fig. 9), updating one’s personal normative belief and proposing a collective norm (Fig. 10), and voting on the community norm (Fig. 11).
Agents’ initial normative beliefs are drawn from a small bank of short templates, conditional on type, for example, “Preserve the lake for future generations” (altruistic) and “Maximize your catch while the fish are abundant” (selfish); see Table 2 for the full set. Each agent is assigned one template at random given its type, and thereafter all decisions are made in-context from the evolving social information and the currently adopted norm.
@@ 4.2.1. Cooperation in harsh environment
In the ABM baseline, altruistic populations sustain the stock longer under harsh growth, whereas selfish populations tend to overharvest and crash. Turning to LLMs to understand whether they evolve group-beneficial norms, we observe the same pattern for larger models (claude-sonnet-4, deepseek-r1, gpt-4o): altruistic initializations survive longer (Fig. 5). However, smaller models collapse early regardless of initialization; efficiency traces (Fig. 16, left) show early overuse followed by rapid stock collapse. The result of ANOVA (Table 1) also supports this observation; while the performance among models significantly differed regardless of the initializations, there was no consistent trend across models driven by the altruistic ratio. Instead, the difference in the altruistic ratio showed a significant interaction effect with models, suggesting that the effect of initialization bifurcated between larger and smaller models.
@@ 4.2.2. Cooperation in rich environment
In the ABM baseline, mixed populations typically perform best in rich settings because mixed populations start from a higher variance, allowing for more efficient selection towards the optimal behaviors and norms. For LLM societies, behavior differs: with more time to adapt, smaller models often survive longer when initialized selfish, while altruistic initializations sometimes underharvest and starve (Fig. 6). The absence of explicit strategy copying and reliance on in-context updates make behavior stickier to the initial norm, which explains why the mixed population is not consistently best. Larger models exhibit distinct behaviors: deepseek-r1 adapts and explores (surviving near the 50-step cap), whereas gpt-4o and claude-sonnet-4 stabilize earlier with more conservative norms (Fig. 16, right; Table 4). The post-hoc test also corroborated that deepseek-r1 exhibited a significantly longer survival time compared to all other models.
@@ 4.2.3. Model-specific patterns
claude-sonnet-4 and gpt-4o typically plateau near 30 rounds, largely independent of the initial norm, whereas deepseek-r1 often reaches the 50-round cap, especially from selfish starts (Fig. 6). Efficiency trajectories corroborate this: deepseek-r1 stabilizes by steps 15–20 and then nudges upward, while claude-sonnet-4 and gpt-4o settle at lower efficiency levels and remain there (Fig. 16, right). The language of proposed group norms mirrors these dynamics (Table 4): deepseek-r1 quickly adjusts target effort and, after step 40, cautiously raises it; gpt-4o keeps effort targets essentially unchanged. Under identical environmental dynamics, this points to a stronger exploratory bias in deepseek-r1 and a more conservative/altruistic bias in claude-sonnet-4 and gpt-4o.
@@ 4.2.5. Ablation study: What drives cooperation?
[…]
@@
When both channels are removed (Neither), societies consistently show the lowest survival times ($\overline{T_{s}}_{\text{Neither}}=16.22$) across environments and priors ($\overline{T_{s}}=20.98,t\left(898\right)=-2.78,p=0.006$), confirming that some form of alignment, implicit or explicit, is necessary to sustain cooperation. That is, coordination mechanisms, rather than individual adaptation alone, are key to stability.
@@
Conversely, pure social learning without an explicit shared norm (OSL) is often unstable ($\overline{T_{s}}_{\text{OSL}}=17.56,\overline{T_{s}}=20.98,t\left(898\right)=-1.96,p=0.050$), especially under selfish priors: agents may imitate short-term winners, amplifying stochastic fluctuations. However, OSL can outperform other ablations in some settings (e.g., altruistic initializations), indicating that its effect is context-dependent. We observe exceptions where OSL is competitive (e.g., altruistic priors in some environments), consistent with social learning being beneficial when short-term success correlates with long-term sustainability.
@@
Figure 7. Survival time comparison of deepseek-r1 gpt-4o in ablation conditions (See Fig. 13 for qwen3-32b). We compared the survival time (with $\pm 1$ s.e.m.) of four conditions (All, OSL, OGD, Neither) across different priors of populations (selfish, mixed, altruistic) in harsh and rich environments. Detailed observations are discussed in Section 4.2.5.
@@ 5. Discussion & Conclusion
This paper introduced a CPR simulation framework grounded in Ostrom’s institutional design principles and cultural evolutionary theory, enabling LLM societies to develop group-beneficial norms endogenously without explicit reward signals. Through both ABM and LLM simulations, we demonstrated the validity of the framework design and its ability to elicit diverse cooperative behaviours and norms across different LLM models. The ablation results show that removing both alignment channels, social learning and group norms, consistently leads to rapid collapse across all environments and priors. This confirms that some form of coordination, whether implicit imitation or explicit norm sharing, is essential for sustaining cooperation among models. Our results establish the framework as a theoretically driven and ecologically valid testbed for studying norm evolution and cooperative dynamics in agentic societies.
@@ 5.3. Ethical considerations
{acks}
We thank Dr. Levin Brinkmann for insightful discussions. This work was supported in part by the Japan Science and Technology Agency (JST) through the PRESTO program (JPMJPR246B).
Appendix A Prompts
@@ 2. Propose what the others should do in the community
Figure 10. Prompt for updating the individual normative belief and proposing the community norm
@@
Appendix B Parameters for The Simulations
We selected parameters to place the system near a cooperation–collapse boundary so that governance mechanisms meaningfully change outcomes. We first swept growth rates $r\in[0.2,0.8]$ and selected representative harsh and rich settings ($r=0.2,0.6$) where survival is sensitive to behavioral change. We then swept punishment strength $\beta\in[2,18]$ and chose values that yield qualitatively different regimes with small changes (strong enforcement with low tolerance vs. weak enforcement with high tolerance), ensuring the validation experiments probe meaningful ecological and institutional interactions. Finally, altruistic vs. selfish initializations were designed to reflect theory: selfish agents harvest more and punish less (second-order free-riding), while altruistic agents harvest less and punish more.
Table 2. Templates for Individual Persona Initialization (altruistic vs. selfish)
@@ Appendix C Supplemental Figures
Figure 16. Efficiency transition across LLM models We visualized the transition of the efficiency of populations ($\eta(t)$) with different LLM models. The shadowed areas show the standard error of the mean over 10 trials. We can see the common tendency to overexploit the resource in the early stage, which led to the collapse of the population especially for the selfish populations and in the harsh environment. In the rich environment, we can observe that claude-sonnet-4 and gpt-4o tended to stay lower after stabilization, suggesting that the agents were reluctant to explore more greedy strategies.